OpenStudy (anonymous):

This person needs help!! Please!! ASAP What is the solution to the equation? 4x – 6 = –3x + 8

OpenStudy (anonymous):

14 or 2

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Well, first you need to isolate the variable. Move all x terms to one side and all others to another side.

OpenStudy (anonymous):

could you show me?

OpenStudy (anonymous):

Doing this gives you 7x = 14. Divide, and x = 2.

OpenStudy (anonymous):

4x+3x=7x 8+6=14 14/7 =2

OpenStudy (anonymous):

Thank you all!!